a.            An economy produces three goods: organ, apple and organ juice. It is assumed that half of the organ are bought and consumed as final good; the other half is used to produce organ juice.                                                                                                 (2 Marks)

In 2019, 20 KGs of organ are produced at Rs.50/kg, 40 KGs of apple are produced at 100/kg, and 25 Glasses of organ juice are produced at Rs.200/glass.

In 2020, 40 KGs of organ are produced at Rs.80/kg, 60 KGs of apple are produced at 110/kg, and 30 glasses of organ juice are produced at Rs.195/glass.

 

  Find the following:

 

i.               Assuming 2019 is the base year; calculate the real and nominal GDP in 2019 and

2020.


b.    Analyze how much Pakistan’s GDP and each of its components is affected by the following transactions? Explain your answers                                                    (3 Marks). 


i.                Atiya spends Rs.1000 to purchase dinner at the finest restaurant in Karachi

ii.              Your mother spend weekend cleaning your house

iii.            Dalda Banspati Limited Pakistan builds Rs.50 million worth of cooking oil, but consumers only buy Rs. 45 million worth of them.



1
Expert's answer
2021-01-27T07:30:58-0500

a) i) Nominal GDP in 2019 "=20\times50+40\times100+25\times200=Rs10000"

Nominal GDP in 2020 "=40\times80+60\times110+30\times195=Rs=15650"

Real GDP in 2019 will be equal to it's nominal GDP because 2019 is the base year. So, real GDP in 2019 will be = Rs 10000

Real GDP in 2020 "=" Quantities of 2020 "\times" Price of 2019 "=40\times50+60\times100+30\times200=Rs14000"


b) i) If Atiya spends Rs. 1000 to purchase dinner at the finest restaurant in Karachi, then this transaction will increase the consumption component by Rs. 1000.

ii) The mother spending weekend cleaning the house will not make any changes in any of the components of the GDP.

iii) In this case, the GDP will increase by the total amount of production, which is Rs. 50 million. This is because GDP is the fin value of the goods and services produced within the boundary of the country. Out of this Rs. 50 million, Rs. 45 million will be increased in the consumption component and Rs. 5 million will be increased in the investment component.